### v1

# OptiChat: Bridging Optimization Models and Practitioners with Large Language Models

Abstract
Optimization models have been applied to solve a wide variety of decision-making problems. These models are usually developed by optimization experts but are used by practitioners without optimization expertise in various application domains. As a result, practitioners often struggle to interact with and draw useful conclusions from optimization models independently. To fill this gap, we introduce OptiChat, a natural language dialogue system designed to help practitioners interpret model formulation, diagnose infeasibility, analyze sensitivity, retrieve information, evaluate modifications, and provide counterfactual explanations. By augmenting large language models (LLMs) with functional calls and code generation tailored for optimization models, we enable seamless interaction and minimize the risk of hallucinations in OptiChat. We develop a new dataset to evaluate OptiChat’s performance in explaining optimization models. Experiments demonstrate that OptiChat effectively bridges the gap between optimization models and practitioners, delivering autonomous, accurate, and instant responses.
1 Introduction
Significant progress in large language models (LLMs) has been witnessed in recent years [1, 2], with applications emerging in chemistry [3, 4, 5, 6], biology [7, 8], healthcare [9, 10], finance [11, 12, 13], manufacturing [14, 15], supply chain management [16], and optimization modeling [17, 18, 19, 20, 21, 22]. This widespread applicability highlights the capability of LLMs to comprehend and process natural language across diverse domains. Leveraging the versatility of LLMs, a recent study introduced an interactive dialogue system, TalktoModel [23], designed to assist practitioners in understanding complex machine learning models from various application domains by integrating LLMs with explainable AI (XAI) techniques [24, 25, 26, 27]. Case studies demonstrated that the interactive dialogue system of TalktoModel significantly enhanced healthcare workers’ understanding of an ML-based disease prediction model.
In addition to machine learning models, optimization models are another type of complex model widely used across various fields, such as engineering, economics, healthcare, and manufacturing [28]. These models formulate real-world decision-making problems into optimizing an objective under a set of constraints (e.g., budgets and operational limits), where each decision variable has a well-defined physical meaning. For example, a supply chain optimization model aims at determining the most economically efficient production levels and product distribution while meeting customer demands. While optimization models are built on structured and interpretable mathematical formulations that are accessible to experts, practitioners face significant challenges in understanding the abstract formulations and often perceive them as black boxes. This challenge becomes even more pronounced when the solutions suggested by these models deviate from familiar heuristics, leaving practitioners uncertain whether to trust them. To address these challenges, the optimization community has also developed explanation techniques analogous to those in XAI. Since the 1980s, the community has developed expert systems [29, 30], interpretable decision rules [31, 32, 33], argumentation-based methods [34, 35, 36], and counterfactual explanations [37].
However, the effective use of these explanation techniques still requires substantial optimization expertise. It can still be challenging for the end users of the optimization models, such as logistics coordinators, to independently comprehend the model’s outcomes, reconcile results that conflict with their own experiences, or conduct further analysis. This limitation imposes a significant burden on the optimization experts tasked with communicating the results to practitioners, slows decision-making processes, and prohibits the dissemination of optimization models in areas where optimization experts are inaccessible.
Figure 1: Overview of OptiChat. Inputs in blue, outputs in green, and tools in grey. (a) User interface. During a conversation, the solution-agnostic query is answered directly through the LLM’s in-context reasoning. The solution-specific query is addressed by interacting with several optimization tools and the model. (b) Backend system. The model code is preprocessed into a natural language description, both of which are accessible to the LLM and the optimization tools. The optimization tools include an optimization solver, an algebraic modeling language, predefined functions, and code generation.
To address this limitation, a natural approach is to leverage LLMs for explaining optimization models, given their versatility in natural language-based tasks. However, the hallucination of existing LLMs is concerning because it can produce unfounded or inaccurate explanations [38]. One solution proposed by Li et al. [16] is to rely entirely on code generation to handle all queries for a supply chain optimization model. However, the prompts developed in their approach are specifically tailored to the supply chain optimization models, which may lead to inaccurate code for general optimization models with complex queries. Alternatively, our prior work [39] augments LLMs with predefined functions, but this system is specifically designed for diagnosing infeasible optimization models. In addition, there are no open-source datasets for testing the natural language systems in explaining optimization models.
Figure 2: Illustrative examples of five different types of queries, the proposed actions, and answers. The proposed actions represent the strategies that generate the relevant explanations, implemented in OptiChat through predefined functions or code generation. The answers are concise summaries of the responses produced by OptiChat.
In view of these challenges, we introduce OptiChat, an LLM-assisted system designed to not only interpret optimization models in natural language but also provide post hoc explanations through interactive dialogues after the optimization model is solved. Common types of user queries can be classified and systematically addressed by OptiChat. To achieve this, we combine various predefined functions and code generation to draw faithful and trustworthy conclusions from optimization models. We also curate a comprehensive dataset to evaluate OptiChat, supporting further research in this area.
As illustrated in Figure 1, the input to OptiChat is a well-written code script corresponding to an optimization model in the Pyomo/Python[40] algebraic modeling language, developed by an optimization expert. OptiChat initially processes the code to generate a coherent and easy-to-understand description to assist practitioners in understanding the problem context. After that, practitioners can ask queries of interest in an interactive dialogue. We survey the queries gathered from practitioners and classify them into five categories: diagnosing, retrieval, sensitivity, what-if, and why-not queries (see illustrative examples in Figure 2). Diagnosing queries focus on fixing infeasibility issues. Retrieval queries seek to extract relevant model data. Sensitivity queries measure the impact of parameter uncertainties on the optimal objective value. What-if queries evaluate significant changes in input parameters specified by users. Lastly, why-not queries investigate the counterfactual scenarios suggested by users. These are termed as solution-specific queries, meaning OptiChat must derive a solution from the model to provide an accurate explanation. Underneath, explanation strategies for each query category are either implemented as predefined functions or generated as code by the LLM, which are then executed using an optimization solver, such as Gurobi. These two approaches are tailored to achieve high success rates in answering queries. Besides the solution-specific queries, the users can also seek qualitative explanations from the LLM. These queries, termed as solution-agnostic queries, are contextualized within the LLM’s memory using the input optimization model and the chat history to ensure the quality of the answers.
2 Results and Discussion
In this section, the effectiveness of OptiChat is demonstrated in terms of model descriptions and query responses. First, we emphasize the efficiency of OptiChat in generating autonomous model descriptions compared to consulting with optimization experts. Next, we measure the success rates of responses for each query type, providing a quantitative evaluation of OptiChat’s accuracy. We then present a qualitative showcase of query responses, illustrating how OptiChat assists practitioners in real-world settings.
We test OptiChat on 24 optimization models written in the Pyomo/Python framework. To demonstrate the versatility of LLMs, the 24 models span various contexts, including supply chain, manufacturing and production, petroleum refinery, industrial scheduling, chemical and process system engineering, transportation, etc. The 24 models are adapted from the GAMS Model Library [^1] https://www.gams.com/latest/gamslib_ml/libhtml/index.html, sourced from the Pyomo Cookbook by the University of Notre Dame [^2] https://jckantor.github.io/ND-Pyomo-Cookbook and a public GitHub tutorial [^3] https://github.com/hdavid16/RTN-Demo, or adapted from an optimization textbook [28]. Infeasible variants of these models are created by adjusting model parameters or introducing additional constraints. An infeasible variant does not refer to code bugs that prohibit models from being executed but rather to an optimization model in which no solution exists to satisfy all constraints simultaneously. This usually occurs because of overly restrictive constraints and inaccuracies in the underlying data. For example, in an aggressive overselling scenario, an aircraft may lack sufficient seats to accommodate all ticketed passengers.
OptiChat aims to make optimization models more accessible to a wider audience by enabling seamless interaction between practitioners and the underlying models, thereby reducing the time experts must spend communicating with practitioners. For our experiments, we recruited 29 experts, including graduate students and postdoctoral researchers, each with at least one year of experience in optimization theory or modeling. These experts were tasked with drafting model descriptions for practitioners and answering follow-up queries. We will show that OptiChat can emulate these experts by producing general model descriptions and responding to user queries in much shorter timeframes.
2.1 Model Description
As practitioners cannot independently interpret optimization models, they must seek assistance from either experts or OptiChat to understand the problem contexts. In this study, the expert participants were required to write detailed model descriptions based on the Pyomo scripts of the optimization models. If a model was found to be infeasible, the expert participants were also responsible for diagnosing the source of the infeasibility. Following these tasks, the experts were surveyed to provide time estimates for completing them. Given the differences in expertise among the participants and the varying complexity of the assigned models, we report the most frequently selected time range in the survey.
On the other hand, OptiChat automatically interprets optimization models in natural language without involving optimization experts, which significantly shortens the time. More importantly, both the experts and OptiChat are augmented with tools to isolate the irreducible infeasible subset (IIS, see details in Section 4.2) for characterizing infeasible optimization models. By isolating the conflicting constraints, optimization experts can gain valuable insight from the IIS analysis. However, it is still time-consuming for them to identify root causes and take corrective actions. In contrast, OptiChat automates the troubleshooting process along with a model description within one minute, as shown in Table 1. Empirically, the quality of OptiChat’s description is observed to be comparable to those provided by experts, which can be attributed to the versatility of LLMs in different problem contexts.
Table 1: Model Description Results
OptiChat Expert
Model Time (min) Time (min)
Feasible $<1$ $18-40$
Infeasible $<1$ $23-55$
[…]
2.2 Query Response Assessment
After reading the autonomous model description, practitioners gain an understanding of the problem context and are prepared to interact with OptiChat by asking queries. Solution-specific queries are addressed in two steps. First, OptiChat either invokes predefined functions or generates code to draw accurate conclusions from the optimization models. Second, OptiChat explains these conclusions in natural language to guarantee their interpretability to practitioners.
2.2.1   Query Dataset Development
We curate a comprehensive dataset of 172 question-answer pairs to evaluate the performance of OptiChat in the interactive dialogue, with each pair developed in the context of a particular optimization model. We classify these questions into 5 categories: (1) diagnosing; (2) retrieval; (3) sensitivity; (4) what-if; (5) why-not queries as exemplified in Figure 2.
Diagnosing Query. Model infeasibility indicates that no decision can be made to simultaneously satisfy all constraints. For instance, it may be impossible to produce a chemical at a 99% concentration with the current resources available in a chemical plant. Restoring feasibility can be approached in various ways, but not all are actionable in the real world. To restore feasibility, increasing the conversion rate of an existing chemical reaction is generally impractical while negotiating with business partners to lower product purity requirements is more actionable. Practitioners seek strategies that align with their operational priorities. To address this need, we develop restoration strategies by introducing slack variables into the original models, as formulated in (2) in the Methods section. The optimal solution found in (2) represents the minimal adjustments needed to restore feasibility.
Retrieval Query. An optimization model consists of decision variables, parameters data, constraints, and an objective function. Optimization models developed for practical use are often large in scale, with some components indexed over large sets. For example, a variable that represents the assignment decision of an aircraft can be indexed over hundreds of routes. When practitioners need to review specific data or optimal decisions from the model, it is inefficient to retrieve this information. OptiChat facilitates real-time access to specific model information through natural language.
Sensitivity Query. Many optimization models are constructed with incomplete knowledge of problem parameters. For example, electricity prices and customer demands often fluctuate in the market over time and cannot be fixed within the model. The values of these parameters can be predicted using historical data and updated on a rolling basis. It is important for practitioners to evaluate how changes in problem parameters impact the optimal objective value, undertake risk assessment, and devise appropriate management strategies. We propose to analyze the sensitivity by leveraging the well-established duality theory and interpreting the optimal dual solution in (3) in Section 4.2. The optimal dual solution can be viewed as the gradient of the optimal objective value with respect to the input parameters in the vicinity where the optimal dual solution is drawn. In other words, it precisely reflects the sensitivity of the optimal objective value.
What-if Query. Although sensitivity analysis can evaluate local changes in parameters, it is not a valid methodology to evaluate larger changes. This limitation arises because the optimal dual solution in (3) in Section 4.2 is not constant but depends on parameters. After large parameter perturbations, the dual solution previously found in the original problem no longer accurately reflects the change in optimal objective value. These significant perturbations often occur when a new policy is being established, or the industry is evaluating a new business strategy. In these cases, OptiChat systematically identifies the extent and type of modification from the user’s query and updates the model to (4) accordingly. Queries such as “increase the labor force to 35 people” and “the customer orders are cut by a third” can generally be handled.
Why-not Query Mathematical optimization is a rigorous methodology that searches for the optimal solution. In contrast, practitioners often rely on their experience to make decisions. As a result, it can be challenging to convince business managers or stakeholders to accept the optimal solutions given by an optimization model, especially when these solutions are against their intuitions. To handle the “why-not” queries, OptiChat treats the scenario proposed by users as a counterfactual and implements it into the original model as a set of constraints. For instance, consider a truck driver who questions why a specific route—one that seems efficient to them—is not chosen by the optimization model. To investigate this, the optimization expert can modify the original model by adding constraints that force the selection of the driver’s preferred route. Solving this modified model produces a new solution, but likely one that leads to higher costs, longer delivery times, or other disadvantages. By comparing these results with the original solution, the expert can demonstrate the trade-offs and show why their proposed route is not optimal. This evidence-based explanation helps build trust in the model’s recommendations. Answering the “why-not” query relies on code generated by the LLM since the counterfactual can be any arbitrary fact suggested by the practitioners, for which predefined functions do not suffice.
Ground Truth Answers To quantitatively evaluate the accuracy of answers, we establish ground-truth answers for each of the 172 questions. The dataset outputs for why-not queries that rely on code generation are natural language-based answers drafted by experts. For the other queries that rely on predefined function calls, the dataset outputs are in JSON format, comprising the function name to invoke and its arguments, such as component names and component indexes.
2.2.2   Quantitative Assessment of the Query Responses
We use the success rate of the query responses as a metric to evaluate OptiChat’s accuracy. OptiChat answers the why-not queries through code generation. The success is verified by manually checking if the execution results of the code generated by LLMs match the expert-drafted answers. For the remaining queries, OptiChat first invokes predefined functions and then provides more natural explanations. The success can be automatically determined by whether the function name and arguments recognized by LLMs match the ground truth. This is because we empirically observe that the LLM does not hallucinate when explaining based on function outputs, as long as the correct function and arguments are chosen.
The quantitative results, as presented in Table 2, demonstrate a $75.7\%$ success rate for why-not queries and above $80\%$ success rate for the remaining queries. OptiChat exhibits a significant advantage in response time, consistently addressing all query types in under one minute. In contrast, experts require one to two orders of magnitude more time on average.
The high accuracy of OptiChat is attributed to two key factors. First, we develop a multi-agent framework to guide LLMs in generating the code for counterfactual constraints. The prompt is specifically tailored for why-not queries rather than general queries, guiding the LLM to generate a minimal amount of code under a narrowed task scope. Second, we incorporate various predefined functions to prevent LLMs from developing explanatory techniques from scratch for other queries. This approach is more robust than code generation, as demonstrated by the higher success rates in the first four queries in Table 2. More importantly, code generation is currently unable to develop more advanced explanatory techniques necessary for diagnosing queries and sensitivity queries. It is observed that code generation only produces heuristic techniques to tackle these queries, which results in suboptimal conclusions. Therefore, predefined functions are indispensable at the current stage.
Nevertheless, the LLM may occasionally fail to match a query to the appropriate function or generate code deviating from users’ purposes. The ambiguity of natural language can also mislead the LLM, which generates function arguments or code that are not executable. Practitioners may describe the same component differently, yet only one executable representation exists in the model. This problem is exacerbated when the component is indexed in multiple dimensions. They might specify only a subset of dimensions, list dimensions in an incorrect order, or omit some entirely. For example, pc["max", :] represents the maximum production capacities for all facilities, but practitioners might refer to it orally as max output limits, often omitting the phrase for all facilities. These failures can be attributed to the under-representation of this specialized area and scarcity of optimization-related data during LLM training.
Table 2: Quantitative Results of Query Responses
OptiChat OptiChat Expert
Query Accuracy (%) Time (min) Time (min)
Diagnosing $87.2$ $0.3$ $13.1$
Retrieval $82.0$ $0.2$ $7.8$
Sensitivity $94.4$ $0.3$ $9.3$
What-if $84.6$ $0.3$ $8.5$
Why-not $75.7$ $0.5$ $9.7$
[…]
Figure 3: A conversation with OptiChat. a, OptiChat helps practitioners identify sensitive parameters
Figure 4: A conversation with OptiChat. b, OptiChat helps practitioners embrace unanticipated decisions
2.2.3   Qualitative Assessment of Query Responses
We showcase some insightful answers produced by OptiChat. The models used for the demonstration are built for supply chain management, which decides the optimal production, storage, and transportation of goods across production facilities, distribution centers, and markets. The maximum capacities of the normal and the overtime production often fluctuate due to factors such as raw material availability and labor engagement. Consider a case where a manufacturer is informed about an increase in raw materials and plans to expand maximum production capacities at certain facilities. As suggested in Figure 3, the optimal profit is highly sensitive to the maximum overtime production capacity at the third facility, even though overtime production is commonly perceived as costly. Consequently, during the strategy development phase, it is crucial to prioritize the third facility over others and produce these additional items at the overtime stage.
The second model takes into account recovery centers that manage re-manufacturing and recovery activities. This model suggests building two such recovery centers, despite the substantial investment cost typically associated with each. Suppose a situation where business managers express concerns regarding the financial burden and are inclined to construct fewer recovery centers. To reassure them, OptiChat justifies this decision by demonstrating that no feasible solution exists if forcing the number of recovery centers no greater than one in Figure 4. Therefore, business managers can be convinced to deprecate outdated practices and appreciate the construction of recovery centers.
3 Conclusions
In this paper, we propose OptiChat to address the common challenge that end users of optimization models are often not optimization experts. By leveraging the recent advancements in LLMs, OptiChat bypasses the need for inefficient back-and-forth coordination with optimization experts. Through straightforward natural language conversations with models, the practitioners can benefit from autonomous and instant responses. The integration of optimization tools with LLMs is demonstrated to be essential for providing a richer and more reliable context for analysis, rather than relying entirely on natural language processing to generate plausible explanations.
Although OptiChat, at its current status, does not achieve perfect accuracy in executing tools, it is anticipated that this issue can be further mitigated by foreseeable advancements in LLM technology and targeted fine-tuning for optimization. On the other hand, the modular nature of OptiChat facilitates easy extension of its functionalities as new explanation techniques emerge. Developing alternative techniques to expand OptiChat’s toolsets is of particular interest. We believe that OptiChat, along with the new dataset and the insights into query classification, could serve as a catalyst in this understudied area and draw more attention from the optimization and the machine learning community.
4 Methods
Figure 5: Multi-agent framework of OptiChat. (a) The Pyomo model code is interpreted by the illustrator agent. Solution-agnostic queries are addressed by the explainer agent directly. Solution-specific queries are handled by a team of engineering agents, followed by the explainer agent. (b) The engineering agents include the reminder, operator, programmer and evaluator agents.
OptiChat is composed of a user interface, a structured sequence of LLM-based agents, an optimization solver, a modeling software, and several custom-built functionalities. In this work, OptiChat utilizes GPT-4 [1] for the LLM-based agents, Gurobi [41] as the optimization solver, and Pyomo [40] as the algebraic modeling language. This section first details the multi-agent framework, outlining the role and sub-task of each agent. Next, we present relevant concepts and mechanisms behind the implementation of OptiChat’s functionalities, connecting them with the proposed explanation strategies.
4.1 Multi-agent Framework
Figure 5 depicts the workflow of the multi-agent framework. Before the interactive dialogue, OptiChat begins with the illustrator agent, which preprocesses the Pyomo optimization model uploaded by users. During the conversation, the coordinator distributes queries to either the explainer or the team of engineering agents. The engineering team operates in a structured order, comprising the following agents: (1) the operator, who controls all predefined functions; (2) the programmer and evaluator agents, responsible for generating code for tasks that cannot be addressed by the predefined functions; and (3) the reminder agent, who provides supplementary information to enhance the accuracy of the other two.
Illustrator This agent extracts sets, parameters, variables, constraints, and objectives from the Pyomo optimization model, labeling each component with a natural language description. The simplified model information, along with other model attributes, is stored for future interactions. If the model is infeasible, the irreducible infeasible subsets (detailed in Section 4.2) are saved as well. Since users may use different terms or notations for specific components, this preprocessing enables the LLMs to accurately identify the referenced component by looking up a table of component symbols defined in the model and their corresponding physical meanings. In addition, OptiChat can access any desired model attributes and interact more effectively with optimization tools. At the user interface, the agent concisely and coherently illustrates the model using natural language rather than overwhelming the user with a complete list of component descriptions and technical model attributes.
Coordinator Each conversation begins and ends with the coordinator agent, who classifies user queries as solution-agnostic or solution-specific. Solution-agnostic queries are solely addressed by the explainer agent through in-context reasoning, while solution-specific queries require technical feedback from agents in the engineering team before being forwarded to the explainer agent.
Explainer The explainer agent is responsible for conveying any technical information to practitioners in an understandable manner.
Operator We implement four predefined functions in the operator agent to address the diagnosing, retrieval, sensitivity, and what-if queries in Figure 2. The LLMs are only responsible for selecting the appropriate function and arguments, after which the operator invokes the corresponding tool to generate the solution. The fundamentals of these predefined functions are introduced in Section 4.2.
Reminder The reminder agent serves as the starting point for every solution-specific query, which helps LLMs more accurately determine error-prone function arguments, such as component indexes. For function calls, LLMs find it relatively easier to identify the appropriate function name through few-shot demonstrations. Additionally, with the component table obtained during preprocessing, LLMs can easily identify the component name of interest. The reminder agent identifies these simpler arguments, generating syntax guidance tailored to the specific function and component, which has been empirically shown to improve the identification accuracy.
Programmer And Evaluator The development of the programmer and evaluator agents is inspired by recent works on applying code generation to optimization [16, 18]. The code generation capability of LLMs has been widely explored and demonstrated impressive results across various tasks [42]. These agents are designed to address the why-not queries in Figure 2, but also potentially handle unexpected queries beyond the reach of predefined functions. The programmer and evaluator are involved in a loop with limited iterations. The evaluator first executes the code generated by the programmer, then reviews the terminal outputs and error messages if available. Throughout this process, the code is automatically refined until a bug-free and comprehensive solution is produced. The prompts for why-not queries are designed to guide the LLM in generating additional constraints from the user’s counterfactual query, thereby narrowing the LLM’s task scope and reducing its code output.
4.2 Optimization Fundamentals
Most practical applications in optimization can be formulated as (mixed-integer) linear programs (MILP/LP):
$\min_{\mathbf{x}}$ $\mathbf{c}^{\top}\mathbf{x}$ (1)
subject to $\mathbf{A}\mathbf{x}$ $\leq\mathbf{b}\text{,}$
$\mathbf{x}$ $\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}\text{.}$
where decisions to make are denoted as $\mathbf{x}$ that can consist of both integer and continuous variables; $\mathbf{c}$ represents the cost coefficients; the problem is subject to linear constraints $\mathbf{A}\mathbf{x}\leq\mathbf{b}$: $\mathbf{A}\in\mathbb{R}^{m\times n}$, $\mathbf{b}\in\mathbb{R}^{m}$.
IIS An irreducible infeasible subset, IIS, is a minimal set of constraints and/or variable bounds within an optimization model that causes infeasibility, defined by two key properties: (i) the IIS itself is infeasible, and (ii) any proper subset of the IIS is feasible. Different algorithms have been developed to isolate IIS, such as deletion filter [43], additive method [44], and hybrid approach [45]. Commercial optimization solvers like CPLEX [46] and Gurobi [41] have implemented variants of these IIS detection algorithms. For a comprehensive review of IIS detection, readers are referred to the monograph [47].
Feasibility Restoration By Adding Slacks Resolving infeasibility can be achieved by either (1) recursively removing constraints from the IIS until the model becomes feasible, or (2) adding slack variables to the optimization problem and allowing for adjustments to input parameters. The second approach is more practical as each constraint reflects an important aspect of the problem and cannot always be removed in practice. In contrast, introducing slack variables into the optimization problem in (1) provides the practitioners with a more concrete and actionable plan for feasibility restoration. Mathematically, the following extended problem is solved
$\min_{\mathbf{x,\delta A^{+},\delta A^{-},\delta b^{+}}}$ $\quad\sum_{(i,j)\in\mathcal{S}_{A}}\big{(}\mathbf{\delta A}_{ij}^{+}+\mathbf{\delta A}_{ij}^{-}\big{)}+\sum_{i\in\mathcal{S}_{b}}\mathbf{\delta b}_{i}^{+}$ (2)
subject to $(\mathbf{A}+\mathbf{\delta A^{+}}-\mathbf{\delta A^{-}})\mathbf{x}\leq\mathbf{b}+\mathbf{\delta b^{+}}\text{,}$
$\mathbf{x}\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p},\quad\mathbf{\delta A^{+},\delta A^{-},\delta b^{+}}\geq 0\text{,}$
where $\mathbf{\delta A^{+}}$, $\mathbf{\delta A^{-}}$, $\mathbf{\delta b^{+}}$ are nonnegative slack variables. Since all the constraints are inequalities, adding nonnegative slacks $\mathbf{\delta b^{+}}$ on the right-hand side relaxes the problem. In contrast, adding the left-hand side slacks, $\mathbf{\delta A^{+}}$ and $\mathbf{\delta A^{-}}$, are less desirable. It is worth noting that because only a subset of parameters can be adjusted in practice, the slack variables differ in dimensionality from the parameters $\mathbf{A}$ and $\mathbf{b}$. The objective is to minimize the total perturbation to the original problem by summing up all the slack variables. In principle, different weights could be assigned to different slack variables, representing the cost of perturbing the corresponding parameters.
Sensitivity Analysis OptiChat performs sensitivity analysis based on well-established duality theory in linear programming (LP) [48] where $\mathbf{x}$ does not contain integer variables. In short, the change in the optimal objective value in response to changes in input parameters can be expressed as a value function
$v(\mathbf{b})=\min_{\mathbf{x}\in\mathcal{X}}\,\mathbf{c}^{\top}\mathbf{x}=\mathbf{y^{*}}^{\top}\mathbf{b}\text{,}$ (3)
where $\mathcal{X}=\{\mathbf{x}:\mathbf{A}\mathbf{x}\leq\mathbf{b}\}$ and $\mathbf{y^{*}}^{\top}$ is the optimal solution to the dual problem of (1). Problem (1) called a primal problem that finds the minimum cost while satisfying all the constraints, whereas the dual problem of (1) aims to seek the tightest lower bound on the optimal primal cost. Denote the original data given in the primal problem as $\mathbf{b^{*}}$, the expression (3) indicates that when the perturbed parameters $\mathbf{b}$ are close to $\mathbf{b^{*}}$, the optimal primal objective value is a linear function of the input parameters with gradient $\mathbf{y^{*}}$. This suggests that the optimal dual solution for the problem with original input parameters can be used to infer the sensitivity of the optimal objective value in the vicinity of $\mathbf{b^{*}}$. The access to dual information in LP is supported by most optimization solvers [41, 46, 49].
Model Modification For what-if queries, the model (1) can be further revised by updating parameters and constraints:
$\min_{\mathbf{x}}$ $\mathbf{\hat{c}}^{\top}\mathbf{x}$ (4)
subject to $\mathbf{\hat{A}}\mathbf{x}$ $\leq\mathbf{\hat{b}}\text{,}$
$\mathbf{x}$ $\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}\text{.}$
In this revision, $\mathbf{\hat{c}}$, $\mathbf{\hat{A}}$ and $\mathbf{\hat{b}}$ are completed by the interaction between LLMs and our predefined function. Modifications will change the feasible region of the problem, generally leading to a different optimal solution.
Counterfactual Explanation A counterfactual explanation addresses “why-not” queries by examining alternative decisions or outcomes that are not currently realized in the optimal solution. For example, one might ask: “Why not choose supplier 1?” in a supply chain model, even though the current optimal solution does not include it. To investigate such a counterfactual scenario, we modify the original model by incorporating an additional constraint that forces the desired alternative to occur. In this example, if $x_{1}$ represents the binary decision to select supplier 1, we add the constraint $x_{1}=1$ to enforce this choice. Once we re-solve the modified problem with this counterfactual constraint in place, we can observe how the objective value and the feasibility of the solution change. For instance, the forced selection of supplier 1 may raise production costs, reduce overall profitability, or even make the problem infeasible. By comparing the new solution to the original one, we gain insight into the trade-offs and underlying reasons that the original optimal solution did not include supplier 1. This counterfactual reasoning thus provides a transparent and actionable explanation of the model’s decision-making process, illustrating which model parameters and constraints are most influential in ruling out the queried alternative.
4.3 Exceptions Management
When practitioners interact with OptiChat, the system is robust in handling exceptions caused by the lack of optimization expertise. In diagnosing queries, adding slack variables to the left-hand side parameters $\mathbf{A}$ will lead to a product of variables between $\mathbf{\delta A^{+}}$, $\mathbf{\delta A^{-}}$ and $\mathbf{x}$ in (2). This results in a non-convex mixed-integer quadratically constrained program (MIQCP), which is often prohibitive to solve. In many situations, left-hand side parameters represent immutable properties. For example, in the constraint $\mathbf{A}\mathbf{x}\leq\mathbf{b}$ where $\mathbf{b}$ is a deadline, $\mathbf{x}$ indicates task status, and $\mathbf{A}$ represents processing times. The processing times are an inherent property of the machines and cannot be changed. When OptiChat detects a request to add slacks to $\mathbf{A}$, it will alert users of this immutability before initiating the MIQCP.
The sensitivity analysis in OptiChat relies on the strong duality of LP. However, practitioners may request sensitivity analysis concerning left-hand side parameters $\mathbf{A}$ in LP models or parameters in MILP models. Unlike the value function of $\mathbf{b}$ in (3), the dependence of the optimal objective value on left-hand side parameters $\mathbf{A}$ cannot be determined based on duality theory. When the optimization model is an MILP, the equality in (3) no longer holds, breaking the connection between optimal objective value and $\mathbf{b}$. In this case, OptiChat will notify them that sensitivity analysis is not supported and suggest providing specific modifications for evaluation if they still wish to address the same queries. This converts a sensitivity query into a what-if query, which can be addressed in a less restrictive manner. Furthermore, when unexpected failures occur in the predefined functions, the programmer and the evaluator agents will be invoked to generate code-based solutions.
5 Code availability
The implementation of OptiChat is released at GitHub: https://github.com/li-group/OptiChat.git.